OpenStudy (anonymous):

Help! Find s25 for the following arithmetic sequence -7,-2,3,8,...

sam (.sam.):

You know how to apply the formula \[\LARGE S_n=\frac{n}{2}[2a_1+(n-1)d]\]?

sam (.sam.):

you gotta find d

sam (.sam.):

and d is just -2-(-7)=5

sam (.sam.):

n will be the number 25, and \(a_1\) is the first term, -7 subsitute in and evaluate

OpenStudy (anonymous):

s(25)=25/2(2(-7)+25-1)5?

OpenStudy (anonymous):

s(25)=625?